# An Image is Worth More Than 16$\times$16 Patches: Exploring Transformers on Individual Pixels

### v1

@@ Abstract
This work does not introduce a new method. Instead, we present an interesting finding that questions the necessity of the inductive bias – locality in modern computer vision architectures. Concretely, we find that vanilla Transformers can operate by directly treating each individual pixel as a token and achieve highly performant results. This is substantially different from the popular design in Vision Transformer, which maintains the inductive bias from ConvNets towards local neighborhoods (e.g. by treating each 16$\times$16 patch as a token). We mainly showcase the effectiveness of pixels-as-tokens across three well-studied tasks in computer vision: supervised learning for object classification, self-supervised learning via masked autoencoding, and image generation with diffusion models. Although directly operating on individual pixels is less computationally practical, we believe the community must be aware of this surprising piece of knowledge when devising the next generation of neural architectures for computer vision.
@@ 1 Introduction
The deep learning revolution can be characterized as a revolution in inductive biases for computer vision. Learning previously occurred on top of manually crafted features, such as those described in [46, 16], which encoded preconceived notions about useful patterns and structures for specific tasks. In contrast, biases in modern features are no longer predetermined but instead shaped by direct learning from data using predefined model architectures. This paradigm shift’s dominance highlights the potential of reducing feature biases to create more versatile and capable systems that excel across a wide range of vision tasks.
Beyond features, model architectures also possess inductive biases. Reducing these biases can facilitate greater unification not only across tasks but also across data modalities. The Transformer architecture [62] serves as a great example. Initially developed to process natural languages, its effectiveness was subsequently demonstrated for images [22], point clouds [66], codes [8], and many other types of data. Notably, compared to its predecessor in vision – ConvNet [42, 28], Vision Transformer (ViT) [22] carries much less image-specific inductive biases. Nonetheless, the initial advantage from such biases is quickly offset by more data (and models that have enough capacity to store patterns within the data), ultimately becoming restrictions preventing ConvNets from scaling further [22].
Of course, ViT is not entirely free of inductive bias. It gets rid of the spatial hierarchy in the ConvNet and models multiple scales in a plain architecture. However, for other inductive biases, the removal is merely half-way through: translation equivariance still exists in its patch projection layer and all the intermediate blocks; and locality – the notion that neighboring pixels are more related than pixels that are far apart – still exists in its ‘patchification’ step (that represents an image with 16$\times$16 patches on a 2D grid) and position embeddings (when they are manually designed). Therefore, a natural question arises: can we completely eliminate either or both of the remaining two inductive biases? Our work aims to answer this question.
inductive bias ConvNet ViT PiT (ours)
@@
Table 1: Major inductive biases in vision architectures. ConvNet [42, 28] has all three – spatial hierarchy, translation equivariance, and locality – with neighboring pixels being more related than pixels farther apart. Vision Transformer (ViT) [22] removes the spatial hierarchy, reduces (but still retains) translation equivariance and locality. We use Pixel Transformer (PiT) to investigate the complete removal of locality by simply applying Transformers on individual pixels. It works surprisingly well, challenging the mainstream belief that locality is a necessity for vision architectures.
@@
Surprisingly, we find locality can indeed be removed. We arrive at this conclusion by directly treating each individual pixel as a token for the Transformer and using position embeddings learned from scratch. In this way, we introduce zero priors about the 2D grid structure of images. Interestingly, instead of training divergence or steep performance degeneration, we obtain better results in quality from the resulting architecture. For easier reference, we name this PiT, short for Pixel Transformer. Note that our goal is not to promote PiT as an approach to replace ViT, but the fact that PiT works so well suggests there is more signals Transformers can capture by viewing images as sets of individual pixels, rather than 16$\times$16 patches. This finding challenges the conventional belief that ‘locality is a fundamental inductive bias for vision tasks’ (see Tab. 1).
@@
In the main paper, we showcase the effectiveness of PiT via three different case studies: (i) supervised learning for object classification, where CIFAR-100 [38] is used for our main experiments thanks to its 32$\times$32 input size, but the observation also generalizes well to ImageNet [19]; (ii) self-supervised learning on CIFAR-100 via standard Masked Autoencoding (MAE) [26] for pre-training, and fine-tuning for classification; and (iii) image generation with diffusion models, where we follow the architecture of Diffusion Transformer (DiT) [51], and study its pixel variant on ImageNet using the latent token space provided by VQGAN [24]. In all three cases, we find PiT exhibits reasonable behaviors, and achieving results better in quality than baselines equipped with the locality inductive bias. This observation is further generalized to fine-grained classification and depth estimation tasks in the appendix.
@@
As a related investigation, we also examine the importance of two locality designs (position embedding and patchification) within the standard ViT architecture on ImageNet. For position embedding, we have three options: sin-cos [62], learned, and none – with sin-cos carrying the locality bias whilst the other two not. To systematically ‘corrupt’ the locality bias in patchification, we perform a pixel permutation before dividing the input into 256-pixel (akin to a 16$\times$16 patch in ViT) tokens. The permutation is fixed across images, and consists of multiple steps that swap a pixel pair within a distance threshold. Our results suggest that patchification imposes a stronger locality prior, and (given PiT) translation equivariance is still indispensable for network designs.
@@
Admittedly, PiT is not as practical as ViT, since treating each pixel as a token will lead to a sequence length much longer than previously adopted for images. This is especially limiting as Self-Attention operations in Transformers demand quadratic computations. In practice, patchification is still arguably the most effective idea that trades quality for efficiency, and locality is still useful. Nevertheless, we believe our investigation delivers a clean, compelling message that locality is not a necessary inductive bias for model design. We believe this finding will be an integral piece of the community knowledge when exploring the next generations of architectures to process images.
2 Related Work
Locality for images.
To the best of our knowledge, most modern vision architectures [28, 50], including those aimed at simplifications of inductive biases [22, 60], still maintain locality in their design. Manually designed visual features before deep learning are also locally biased. For example, SIFT [46] uses a local descriptor to represent each point of interest; HOG [16] normalizes the gradient strengths locally to account for changes in illumination and contrast. Interestingly, with these features, bag-of-words models [12, 41] were popular – analogous to the set-of-pixels explored in our work.
Locality beyond images.
The inductive bias of locality is widely accepted beyond modeling 2D images. For text, a natural language sequence is often pre-processed with ‘tokenizers’ [57, 39], which aggregate the dataset statistics for grouping frequently-occurring adjacent characters into sub-words. Before Transformers, recurrent neural networks [47, 31] are the default architecture for such data, which exploit the temporal connectivity to process sequences step-by-step. For even less structured data (e.g. point clouds [6, 15]), modern networks [52, 66] will resort to various sampling and pooling strategies to increase their sensitivity to the local geometric layout. In graph neural networks [56], nodes with edges are often viewed as being locally connected, and information is propagated through these connections to farther-away nodes. Such a design make them particularly useful for analyzing social networks, molecular structures, etc.
Figure 1: Overview of a Pixel Transformer (PiT), which is used to investigate the role of locality. Given an image, we simply treat it as a set of pixels. Besides, we also employ randomly initialized and learnable position embeddings without any information about 2D structure, therefore removing the remaining locality inductive bias from previous work (e.g., ViT [23]). These representations are then fed into the Transformer which performs set operations in its interleaved Self-Attention and MLP blocks (only showing one of each for clarity). We demonstrate the versatility of PiT with three case studies, covering both discriminative and generative tasks.
Other notable efforts.
  • For ConvNets, relevant attempts have being made to remove locality. Notably, [4] replaces all the spatial convolutional filters with 1$\times$1 filters in a ResNet [28]. It provides more interpretability to understand the decision making process of a ConvNet, but without inter-pixel communications, the resulting network is substantially worse in performance. Our work instead uses Transformers, which are inherently built on set operations, with the Self-Attention mechanism handling all-to-all communications; understandably, we attain better results.
  • Before ViT gained popularity, iGPT [7] was proposed to directly pre-train Transformers on pixels following their success on text [53, 20]. In retrospect, iGPT is a locality-free model for self-supervised next (or masked) pixel prediction. But despite the expensive demonstrations, its performance still falls short compared to simple contrastive pre-training [9] for ImageNet linear classification. Later, ViT [23] re-introduced locality (e.g., via patchification) into the architecture, achieving impressive results on many benchmarks including ImageNet. Since then, the community has moved on with 16$\times$16 patches as the default tokens for images. Even today, it is still unclear whether higher resolution or locality is the key differentiator between the two. Our work closes this understanding gap, pointing to resolution as the enabler for ViT, not locality with systematic analyses.
  • Perceiver [35, 34] is another series of architectures that operate directly on pixels for images. Aimed at being modality-agnostic, Perceiver designs latent Transformers with cross-attention modules to tackle the efficiency issue when the input is high-dimensional. However, this design is not as widely adopted as plain Transformers, which have consistently demonstrated scalability across multiple domains [5, 22]. Through PiT, we show Transformers can indeed work directly with pixels, and given the rapid development of Self-Attention implementations to handle massive sequence length (up to a million[17, 44], efficiency may not be a critical bottleneck even when all the pixels are counted.
  • Our work can also be viewed as exploring sequence length scaling to the extreme. It’s to model individual pixels for images, and to model individual characters for text [36]. Longer sequence (or higher resolution) is generally beneficial, as evidenced in [10, 32, 51]. However, all of them stopped short of reaching the extreme case that completely gets rid of locality.
@@ 3 Inductive Bias of Locality
In this section, we provide in-depth discussions about the inductive bias of locality (or locality) in mainstream architectures. To be precise, locality is the inductive bias that neighboring pixels are more related than pixels farther apart. We cover ConvNets and ViTs.
@@ 3.1 Locality in ConvNets
In a ConvNet, the locality bias is reflected in the receptive fields of the features computed in each layer of the network. Intuitively, receptive fields cover the pixels involved in computing a specific feature, and for ConvNets, these fields are local.
Specifically, ConvNets consist of several layers, each containing convolutional operations using kernels (e.g., 7$\times$7 or 3$\times$3) or pooling operations – both of which are locally biased. For example, the receptive field in the first layer often corresponds to only a small local window. The field is progressively expanded as the network becomes deeper, but the window is still local and centered at the location of the pixel.
@@ 3.2 Locality in Vision Transformers
At the first glance, Transformers are locality-free. This is because the majority of Transformer operations are either global (e.g., Self-Attention), or purely within each individual token (e.g., MLP). However, a closer look will reveal two designs within ViT [22] that can still retain the locality inductive bias: patchification and position embedding.
@@
In ViT, the tokens fed into the Transformer blocks are patches, not pixels. Each patch consists of 16$\times$16 pixels, and becomes the basic unit of operation after the first projection layer. This means the amount of computation imposed within the patch is drastically different from the amount across patches: the information outside the 16$\times$16 neighborhood can only be propagated with Self-Attention, and the information among the 256 pixels are always processed jointly as token. While the receptive field becomes global after the first Self-Attention block, the bias toward local neighborhood is already inducted in the patchification step.
@@
Position embeddings can be learned [22], or manually designed and fixed during training. A natural choice for images is to use a 2D sin-cos embedding [10, 26], which extends from the original 1D one [62]. As sin-cos functions are smooth, they tend to introduce locality biases that nearby tokens are more similar in the embedding space. [^1] While sin-cos functions are also cyclic, it’s easy to verify that the majority of their periods are longer than the typical sequence lengths encountered by ViTs. Other designed variants are also possible and have been explored [22], but all of them can carry information about the 2D grid structure of images, unlike learned position embedding which does not make assumptions about the input.
The locality bias has also been exploited when the position embeddings are interpolated [18, 43]. Through bilinear or bicubic interpolation, spatially close embeddings are used to generate a new embedding of the current position, which also leverages locality as a prior.
Compared to ConvNets, ViTs are designed with much less pronounced bias toward locality. We push this further by completely removing this bias next.
4 Transformers on Pixels
We closely follow the standard Transformer encoder [62] which processes a sequence of tokens. Particularly, we apply the architecture directly on an unordered set of pixels from the input image with learnable position embeddings. This removes the remaining inductive bias of locality in ViT [22], and for reference, we name it Pixel Transformer (PiT, see Fig. 1). Conceptually, PiT can be viewed as a simplified version of ViT, with 1$\times$1 patches instead of 16$\times$16.
@@ 4 Transformers on Pixels
where $Z_{0}$ is the input sequence $X$, $k\in\{1,...,N\}$ indicates $k$-th layer in the Transformer, and $\mathrm{norm}(\cdot)$ is a normalization layer (typically LayerNorm [1]). Both blocks use residual connections [28].
@@
The typical input in computer vision to the network is commonly an image of RGB values, $I\in\mathbb{R}^{H\times W\times 3}$, where $(H,W)$ is the size of the original image. We follow a simple solution and treat $I$ as an unordered set of pixels $(p_{l})_{l=1}^{H\cdot W},p_{l}\in\mathbb{R}^{3}$. Thus, PiT simply projects each pixel into a $d$ dimensional vector via a linear projection layer, $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{d}$, resulting the input set of tokens $X=\left(f(p_{1}),...,f(p_{L})\right)$ with $L=H\cdot W$. We append the sequence with a learnable [cls] token [20]. Additionally, we learn an content-agnostic position embedding for each position. The pixel tokens are then fed into the Transformer to produce the set of representations $Z$.
@@
PiT-T(iny) 12 192 768 12 5.6
PiT-S(mall) 12 384 1536 12 21.8
PiT-B(ase) 12 768 3072 12 86.0
PiT-L(arge) 24 1024 4096 16 303.5
Table 2: Specifications of PiT size variants. In empirical evaluations, we utilize ViT [22] with the same configuration for head-on comparisons.
PiT removes the locality inductive bias and is permutation equivariant at the pixel level. By treating individual pixels directly as tokens, we assume no spatial relationship in the architecture and let the model learn it from data. This is in contrast to the design of the convolution kernel in ConvNets or the patch-based tokenization in ViT [22], which enforces an inductive bias based on the proximity of pixels. In this regard, PiT is more versatile – it can naturally model arbitrarily sized images (no need to be divisible by the stride or patch size), or even generalize to irregular regions [37].
Besides the removal of locality, using each pixel as a separate token has additional benefits. Similar to treating characters as tokens for language, we can greatly reduce the vocabulary size of input tokens to the Transformer. Specifically, given the pixel of three color channels in the range of $[0,255]$, the maximum size of vocabulary is $255^{3}$ (as pixels take discrete integer values); a patch token of size $p{\times}p$ in ViT, however, can lead to a vocabulary size of up to $255^{3{\cdot}p{\cdot}p}$. If modeled in a non-parametric manner, this will heavily suffer from out-of-vocabulary issues.
Of course, PiT also has downsides, with the biggest one being computationally expensive (or even prohibitive) for modeling long sequences. However, given the rapid development of techniques that handle massive sequence lengths for large language models (up to a million[17, 44], it is entirely possible that soon, we can train PiTs on all the pixels directly (e.g., a standard 224$\times$224 crop on ImageNet ‘only’ contains 50,176 pixels). Therefore, the goal of our paper is to empirically verify the effectiveness and potential of PiT at a smaller scale – which we do next, and leave the engineering effort of practical deployment for the future.
5 Experiments for PiT
In this section, we verify the effectiveness of PiT with three case studies: supervised learning, self-supervised learning with MAE [26], image generation with DiT [51]. We use four variants of PiT: Tiny (T), Small (S), Base (B) and Large (L) with the specifications shown in Tab. 2. Unless otherwise specified, we use the ViT [23] variants of the same configuration as our baselines. In sum, our experiments show that PiT can indeed learn strong vision representations with no inductive bias on locality.
model Acc@1 Acc@5

### v2

# An Image is Worth More Than $\times16$ Patches: Exploring Transformers on Individual Pixels

@@ Abstract
This work does not introduce a new method. Instead, we present an interesting finding that questions the necessity of the inductive bias of locality in modern computer vision architectures. Concretely, we find that vanilla Transformers can operate by directly treating each individual pixel as a token and achieve highly performant results. This is substantially different from the popular design in Vision Transformer, which maintains the inductive bias from ConvNets towards local neighborhoods (e.g. by treating each $\times16$ patch as a token). We showcase the effectiveness of pixels-as-tokens across three well-studied computer vision tasks: supervised learning for classification and regression, self-supervised learning via masked autoencoding, and image generation with diffusion models. Although it’s computationally less practical to directly operate on individual pixels, we believe the community must be made aware of this surprising piece of knowledge when devising the next generation of neural network architectures for computer vision.
@@ 1 Introduction
For computer vision, the deep learning revolution can be characterized as a revolution in inductive biases. Learning previously occurred on top of manually crafted features, such as those described in Lowe (2004); Dalal & Triggs (2005), which encoded preconceived notions about useful patterns and structures for specific tasks. In contrast, biases in modern features are no longer predetermined but instead shaped by direct learning from data using predefined model architectures. This paradigm shift’s dominance highlights the potential of reducing feature biases to create more versatile and capable systems that excel across a wide range of vision tasks.
Beyond features, model architectures also possess inductive biases. Reducing these biases can potentially facilitate greater unification not only across tasks but also across data modalities. The Transformer architecture (Vaswani et al., 2017) serves as a great example. Initially developed to process natural languages, its effectiveness was subsequently demonstrated for images (Dosovitskiy et al., 2021), point clouds (Zhao et al., 2021), codes (Chen et al., 2021a), and many other types of data. Notably, compared to its predecessor in vision – ConvNets (LeCun et al., 1989; He et al., 2016), the Vision Transformer (ViT) (Dosovitskiy et al., 2021) carries much less image-specific inductive bias. Nonetheless, the initial advantage from such biases is quickly offset by more data (and models that have enough capacities to store patterns observed within the data), ultimately becoming restrictions that prevent ConvNets from scaling further (Dosovitskiy et al., 2021).
Of course, ViT is not entirely free of inductive bias. It gets rid of the spatial hierarchy in the ConvNet and models multiple scales in a plain architecture. However, for other inductive biases, the removal is merely half-way through: location equivariance still exists in its patch projection layer and all the intermediate blocks; [^1] By ‘location equivariance’, we refer to the adoption of weight-sharing mechanism which ensures that the same weights are applied regardless of spatial locations. and locality – the notion that neighboring pixels are more related than pixels that are farther apart – still exists in its ‘patchification’ step (that represents an image with $\times16$ 2D patches) and position embeddings (when they are manually designed). Therefore, a natural question arises: can we completely eliminate either or both of these two major inductive biases? In this work, we aim to answer this important question.
inductive bias ConvNet ViT our work
@@
Table 1: Major inductive biases for vision. A ConvNet (LeCun et al., 1989; He et al., 2016) has all three: spatial hierarchy, location equivariance, and locality, with neighboring pixels being more related than pixels farther apart. Vision Transformer (ViT) (Dosovitskiy et al., 2021) removes the spatial hierarchy, reduces (but still retains) location equivariance and locality. We investigate the complete removal of locality by simply applying Transformers on individual pixels. It works surprisingly well, challenging the mainstream belief that locality is a necessity for vision architectures.
@@
Surprisingly, we find locality can indeed be removed. We arrive at this conclusion by directly treating each individual pixel as a token for the Transformer and using position embeddings learned from scratch. In this way, we introduce zero priors about the 2D grid structure of images. Interestingly, instead of training divergence or steep performance degeneration, we can obtain better results in quality from this setup. The fact that this naïve, locality-free Transformer works so well suggests there are more signals Transformers can capture by viewing images as sets of individual pixels, rather than $\times16$ patches. This finding challenges the conventional community belief that ‘locality is a fundamental inductive bias for vision tasks’ (see Tab. 1).
@@
We showcase the effectiveness of Transformer on pixels via three different case studies: (i) supervised learning for object classification, where CIFAR-100 (Krizhevsky, 2009) is used for our main experiments thanks to its default $\times32$ input size, but the observation also generalizes well to ImageNet (Deng et al., 2009), fine-grained classification using Oxford-102-Flowers (Nilsback & Zisserman, 2008) and depth estimation via regression using NYU-v2 (Silberman et al., 2012); (ii) self-supervised learning on CIFAR-100 via Masked Autoencoding (MAE) (He et al., 2022) for pre-training, and fine-tuning for classification; and (iii) image generation with diffusion models, where we follow the architecture of the Diffusion Transformer (DiT) (Peebles & Xie, 2023), and study its pixel variant on ImageNet using the latent token space provided by VQGAN (Esser et al., 2021). In all three cases, we find Transformers on pixels exhibit reasonable behaviors, and achieves results better in quality than baselines equipped with the locality inductive bias.
@@
As a related investigation, we also examine the importance of two locality designs (position embedding and patchification) within the standard ViT architecture on ImageNet. For position embedding, we have three options: sin-cos (Vaswani et al., 2017), learned, and none – with sin-cos carrying the locality bias whilst the other two do not. To systematically ‘corrupt’ the locality bias in patchification, we perform a pixel permutation before dividing the input into 256-pixel (akin to a $\times16$ patch in ViT) tokens. The permutation is fixed across images, and consists of multiple steps that each swaps a pixel pair within a distance threshold. Our results suggest that patchification imposes a stronger locality prior, and removing both location equivariance and locality is not yet feasible.
@@
Treating each pixel as a token will lead to a sequence length much longer than commonly used for images. This is especially cumbersome as Self-Attention in Transformers demands quadratic computations. Therefore, the contribution of our work is on the finding, not on proposing a new method. In practice, patchification is still a simple and effective idea that trades quality for efficiency, and locality is still highly useful. Nevertheless, we believe our investigation delivers a clean, compelling message that locality is not a necessary inductive bias for vision. We believe this finding will be integrated into the community knowledge when exploring the next generations of vision architectures.
@@ 2 Inductive Bias of Locality
Here we discuss about the inductive bias of locality in mainstream vision architectures: ConvNets and ViTs. Locality is the notion that neighboring pixels are more related than pixels farther apart.
Figure 1: Transformer on pixels, or $\times1$ ‘patches’, which we use to investigate the role of locality. Given an image, we simply treat it as a set of pixels. Besides, we also use randomly initialized and learnable position embeddings without prior about the 2D image grid, thus removing the remaining locality bias from previous architectures (e.g., ViT (Dosovitskiy et al., 2014)) that operate on non-degenerated patches. Transformers are employed on the top, with interleaved Self-Attention and MLP blocks (only showing one pair for clarity). We showcase the effectiveness of this locality-free architecture through three case studies, spanning both discriminative and generative tasks.
@@ 2.1 Locality in ConvNets
In a ConvNet, locality is reflected in the receptive fields of the features computed in each layer. Intuitively, receptive fields cover all the pixels involved in computing a specific feature, and for ConvNets, the fields are local. Specifically, ConvNets consist of several layers, each having convolutional kernels or pooling operations – both of which are local. The field is progressively expanded as the network stacks more layers, but the window is still locally centered at the location of the pixel.
@@ 2.2 Locality in Vision Transformers
At first glance, Transformers are locality-free. This is because the majority of Transformer operations are either global (e.g., Self-Attention), or purely within each individual token (e.g., MLP). However, a closer look will reveal two designs within ViT (Dosovitskiy et al., 2021) that can still retain the locality inductive bias: patchification and position embedding.
@@
In ViT, the tokens fed into the Transformer blocks are patches, not pixels. Each patch consists of $\times16$ pixels, and becomes the basic unit of operation after the first projection layer. This means the amount of computation imposed within the patch is substantially different from the amount across patches: the information outside the $\times16$ neighborhood can only be propagated with Self-Attention, and the information among the 256 pixels are always processed jointly as token. While the receptive field becomes global after the first Self-Attention block, the bias towards local neighborhoods is already inducted in the patchification step.
@@
Position embeddings can be learned (Dosovitskiy et al., 2021), or manually designed and fixed during training. A natural choice for images is to use a 2D sin-cos embedding (Chen et al., 2021b; He et al., 2022), which extends from the original 1D one (Vaswani et al., 2017). As sin-cos functions are smooth, they tend to introduce locality biases that nearby tokens are more similar in the embedding space. [^2] While sin-cos functions are also cyclic, it’s easy to verify that the majority of their periods are longer than the typical sequence lengths encountered by ViTs. Other designed variants are also possible and have been explored (Dosovitskiy et al., 2021), but all of them can carry information about the 2D grid structure of images, unlike learned ones which do not make any assumption about the input.
The locality bias has also been exploited when the position embeddings are interpolated (Dehghani et al., 2023; Li et al., 2021). Through bilinear or bicubic interpolation, spatially close embeddings are used to generate a new embedding of the current position, which also leverages locality as a prior.
Compared to ConvNets, ViTs are designed with much less pronounced bias toward locality. This paves the way for our investigation to completely remove this bias, described next.
3 Exploring Transformers on Individual Pixels
To study the impact of removing the locality inductive bias, we closely follow the standard Transformer encoder (Vaswani et al., 2017) which processes a sequence of tokens. Particularly, we apply the architecture directly on an unordered set of pixels from an input image with learnable position embeddings. This removes the remaining inductive bias of locality in the ViT (Dosovitskiy et al., 2021) (see Fig. 1). Conceptually, the resulting architecture can be viewed as a simplified version of ViT, with degenerated $\times1$ patches instead of non-degenerated ones (e.g., $\times16$ or $\times2$).
@@ 3 Exploring Transformers on Individual Pixels
where $Z_{0}$ is the input sequence $X$, $k\in\{1,...,N\}$ the layer index, and $\mathrm{norm}(\cdot)$ is normalization (e.g., LayerNorm (Ba et al., 2016)). Both blocks use residual connections (He et al., 2016).
@@
Given an image $I\in\mathbb{R}^{H\times W\times 3}$ of RGB values, we denote $(H,W)$ as the size of the original input. We follow a simple philosophy and treat $I$ as an unordered set of pixels $(p_{l})_{l=1}^{H\cdot W},p_{l}\in\mathbb{R}^{3}$. The first layer simply projects each pixel into a $d$ dimensional vector via linear projection, $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{d}$, resulting in the input set of tokens $X=\left(f(p_{1}),...,f(p_{L})\right)$ with $L=H\cdot W$. We learn a content-agnostic position embedding for each position, and optionally append the sequence with a learnable [cls] token (Devlin et al., 2019). The pixel tokens are then fed into the Transformer to produce the set of representations $Z$.
@@
T(iny) 12 192 768 12 5.6
S(mall) 12 384 1536 12 21.8
B(ase) 12 768 3072 12 86.0
L(arge) 24 1024 4096 16 303.5
Table 2: Specifications of size variants explored. We use ViT (Dosovitskiy et al., 2021) or DiT (Peebles & Xie, 2023) with the same size but non-degenerated patches for head-on comparisons.
Pixel transformer removes the locality inductive bias and is permutation equivariant at the pixel level. By treating individual pixels directly as tokens, we assume no spatial relationship in the architecture and let the model learn structures directly from data. This is in contrast to the design of the convolution kernel in ConvNets or the patch-based tokenization in ViT (Dosovitskiy et al., 2021), which enforces an inductive bias based on the proximity of pixels. In this regard, the resulting model is more versatile – it can naturally model arbitrarily sized images (no need to be divisible by the stride or patch size), or even generalize to irregular regions (Ke et al., 2022).
While our focus is on the study of locality, using each pixel as a separate token has additional benefits. Similar to treating characters as tokens for language, we can greatly reduce the vocabulary size of input tokens to the Transformer. Specifically, given the pixel of RGB channels in the range of $[0,255]$, the maximum size of vocabulary is $256^{3}$ (as pixels take integer values); a patch token of size $p{\times}p$ in ViT, however, can lead to a vocabulary size of up to $256^{3{\cdot}p{\cdot}p}$. If modeled in a non-parametric manner, this will heavily suffer from out-of-vocabulary issues.
Of course, Transformers on pixels are computationally expensive (or even prohibitive). However, given the rapid development of techniques that handle massive sequence lengths for large language models (up to a million(Dao et al., 2022; Liu et al., 2023), it is entirely possible that soon, we can train Transformers on individual pixels directly (e.g., a standard $\times224$ crop on ImageNet ‘only’ contains 50,176 pixels). In this sense, our work is to scientifically verify the effectiveness and potential of locality-free architectures at an affordable scale (which we do next), and leave the engineering effort of scaling for the future.
4 Experiments
We verify the effectiveness of the locality-free modification with three case studies: supervised learning with ViT (Dosovitskiy et al., 2014), self-supervised learning with MAE (He et al., 2022), and image generation with DiT (Peebles & Xie, 2023). We use four size variants: Tiny (T), Small (S), Base (B) and Large (L) with the specifications shown in Tab. 2. Unless otherwise noted, we use Transformers with the same size but non-degenerated patches as our baselines. Our only goal in this section is to show locality-free architectures can still learn strong vision representations.
Acc@1 Acc@5